10. A 1500 kg yellow convertible is travelling south, and a 2000 kg red station wagon is travelling west. If the total momentum is of the system consisting of the two cars is 8000 kg.m/s directed at 30° west of south, what is the speed of each car?

1
Expert's answer
2022-02-10T08:36:31-0500

Answer

The momentum of the two car system can be found from equations:


((m1v1)² + (m₂v₂)² = p²


cos(60) =


((m₂v¹₂)² + (m₂v¹₂)² = p² 1 (m, + m₂) ₂


P


8000kgm/s


2(1500kg + 2000kg)


12₂ = 2(m₂ + m₂)


= 1.14m/s


From the first equation:


1 m₂ V₁ = -√√p² - (m₂1₂) ²


Py = 1500kg √(8000kgm/s)²-(2000kg 1.14m/s)² = 5.11m/s


Answer: v1=5.11m/s, v2= 1.14m/s